Can you take me step by step in helping me with the prime factorization for the numbers 34, 37, 67, 123, and 71? I would love to learn how to do theses so that I may understand when testing time comes around.
Thank you.

Answer by coyote(32)

Pick 2 numbers that multiply to equal the original number, and keep breaking them down until they are prime (no longer divisible by anything other than 1 and itself).
-----
HINT: Prime numbers will ALWAYS be odd because an even number will be divisible by 2.
-----
Examples
34=2*17 (both prime)
37=already prime
67=already prime
123=3*41
71=already prime
----
Here's another hint: For larger numbers, if you add each number together, you can tell if it's divisible by 3.
Example:336...add 6+3+3=12. 12 IS divisible by 3. So prime factorization on 336 is 3*112
112 can go further, so try 2*56
Still can break down, so try 4*14
And more...2*2*2*7
So final answer is the primes of 336 = 3, 2, 2, 2, 2, 7, althought they like to see it in order, so 2, 2, 2, 2, 3, 7
-----
To check, multiply the factors
2*2*2*2*3*7=336
336=336
Hope that helps!